# MNSC-0001-2024.00

### v1

@@ 2 Network Model Definition
This example illustrates how the failure of a pure principal (e.g., $A$ defaults) can propagate through the network in the stochastic process. If $C$ cannot complete its contractual obligations to $E$ because $A$ fails, then $E$ may incur losses, even though it never directly contracted with $A$. Such indirect dependencies are not captured in standard models assuming independent risk, but they are central in our network-aware framework. This further underscores the role of intermediaries like $C$ in amplifying risk: even with moderate idiosyncratic risk levels for $A$ and $B$, the dependency structure means that failures can cascade through the network, elevating systemic risk beyond what node-level scores would suggest.
@@ 2.1 Discussion of Modeling Assumptions
We treat each contractor $i$’s idiosyncratic risk score $r_{i}$ and network sensitivity parameter $\alpha_{i}$ as exogenously specified inputs to the model. In practice, these parameters are inferred from financial health indicators or historical default data, and likely take into account mild network risk indicators (kim2019default). However, regardless of how $r_{i}$ are estimated, our model allows for the direct incorporation of network effects on risk in contractor networks.
@@
Our stochastic process assumes that a contractor’s risk of failure increases linearly based on the impact of their neighbors via $w_{ij}$ (see Eq. 1). This additive structure simplifies both analysis and simulation, but it may fail to capture important nonlinearities in real-world contagion effects. For instance, a contractor may be robust to isolated failures but vulnerable to risks beyond a certain threshold, such as in the threshold contagion model like watts2002simple.
@@
Our work assumes that network-induced risk depends on the proportion of a contractor’s total subcontracted value attributed to each subcontractor, as encoded by the normalized edge weights $w_{ij}$. This formulation reflects a reasonable first-order approximation: risk exposure grows with financial dependence on risky neighbors. Although real-world contracts vary in risk beyond their dollar value, introducing such heterogeneity would significantly complicate the model without materially changing our core theoretical insights.
@@ 3 Mean Field Analysis of Expected Risk
We start off our analysis by considering the marginal expected risk failure probabilities of the stochastic process $\mathbf{X}^{t}$ for each node $i$. We will later see that this corresponds to a modified Friedkin-Johnsen model in the opinion dynamics literature (FriedkinJohnsen), and calculate a closed-form expression for the mean failure probabilities. We also exploit this representation to describe an eigenvector-based centrality measure, assigning scores to each node in the graph corresponding to their risk-based centrality within the contractor network. We close this section by providing a simple monotonicity condition under which the mean failure probabilities for each node increase due to network effects.
@@ 3.1 Expected Failure Probabilities
The operator $(\mathbf{I}-\mathbf{A}\mathbf{W})^{-1}$ admits multiple interpretations across related literatures. In Markov chain theory, it parallels the fundamental matrix of absorbing chains, where each entry gives the expected number of visits to a state prior to absorption (kemeny1969finite). From a graph-theoretic perspective, it plays a role similar to the pseudoinverse of the graph Laplacian, which encodes mean hitting times, commute times, and effective resistances (lovasz1993random, doyle1984random, chung1997spectral). In our context, this highlights that $(\mathbf{I}-\mathbf{A}\mathbf{W})^{-1}$ aggregates contributions from all higher-order paths in the contractor network, providing an interpretable link between local idiosyncratic risks and their amplified global effects.
@@
lemmaSecondNormBounded Let $\mathbf{A}$ be a diagonal matrix with $\mathbf{A}_{ii}=\alpha_{i}\text{ for all }i\in\mathcal{V}$ and let $\mathbf{W}$ be the row-stochastic adjacency matrix of the contractor graph $G$. Then we have $\lVert(\mathbf{A}\mathbf{W})^{2}\rVert<1$ and $(\mathbf{A}\mathbf{W})^{t}\to 0$ as $t\to\infty$.
This bound is useful for showing that the Neumann series of $\mathbf{A}\mathbf{W}$ converges, which we formalize in the following corollary: {restatable}corollaryNeumann Let $\mathbf{A}$ be a diagonal matrix with $\mathbf{A}_{ii}=\alpha_{i}\text{ for all }i\in\mathcal{V}$ and let $\mathbf{W}$ be the row-stochastic adjacency matrix of the contractor graph $G$. Then $(\mathbf{I}-\mathbf{A}\mathbf{W})$ is invertible, and the Neumann series $\sum_{t=0}^{\infty}(\mathbf{A}\mathbf{W})^{t}$ converges to $(\mathbf{I}-\mathbf{A}\mathbf{W})^{{-1}}$.
@@
Relation to FriedkinJohnsen.
Our model is closely related to the social influence model under static conditions introduced by FriedkinJohnsen, which describes a process in which an individual’s opinion at time $t+1$ is represented as a real-valued linear function of both their innate opinion (determined by exogenous variables) and the opinions of their in-neighbors in the previous time step $t$. Let $\mathbf{m}^{t}$ denote the vector of opinions at time $t\geq 0$, where $\mathbf{r}=\mathbf{m}^{0}$ denotes the vector of inherent opinions that each individual holds at time $t=0$. Then opinions evolve according to the following dynamics for all $t\in\mathbb{N}$:
@@
Adapting their model to our context of risk, we interpret a node’s opinion as its probability of failure, which is dependent on failure probabilities of subcontractors according to the transition matrix $\mathbf{W}$ in the same manner that individuals in the social influence model take their neighbors’ opinions into account. Thus in our context we require that “opinions” take values in $[0,1]$. Moreover, to ensure that $\mathbf{m}^{t}$ maintains its interpretation as a vector of probabilities for all $t\in\mathbb{N}$, we also require $\mathbf{W}$ to be row-stochastic and that $\alpha,\gamma>0$ satisfy $\alpha+\gamma=1$. Then we can rewrite Equation 7 as $\mathbf{m}^{t+1}=(1-\alpha)\mathbf{r}+\alpha\mathbf{W}\mathbf{m}^{t}.$
Comparing with our dynamics in Equation 3, the main difference from FriedkinJohnsen is that they assume a common bias parameter $\alpha$, whereas our model allows node-specific bias parameters by replacing $\alpha$ and $\gamma$ with diagonal matrices $\mathbf{A}$ and $(\mathbf{I}-\mathbf{A})$. Note that in contrast to the Friedkin-Johnsen model, we consider a directed network, so to handle nodes $i\in\mathcal{V}$ that lack in- or out-edges we fix their $\alpha_{i}$ as described in Section 2. More broadly, their framework models the deterministic evolution of average opinions, while ours defines a discrete stochastic process over binary outcomes, yielding the full distribution of failures across the network rather than only mean behavior.
@@ 3.2 Identifying the Impact of Network Structure
where $n$ is the total number of nodes in the network. This expression shows that the overall failure risk can be rewritten as a network-adjusted re-weighting of the individual idiosyncratic risk scores $\mathbf{r}$. The term $(\mathbf{I}-\mathbf{A}\mathbf{W})^{-1}$ highlights that risk exposure is not limited to direct neighbors: it aggregates contributions from all paths in the network, with longer paths down-weighted by successive products of exposure probabilities.
@@
This definition is analogous to traditional PageRank in web search algorithms (BRIN1998107, page1999pagerank, kleinberg1999authoritative), where importance is assigned based on structural connectivity. Here, however, the risk-based centrality captures how individual nodes influence the system-wide failure probability. It accounts for both direct contributions from individual risk levels (through $\mathbf{A}$) and the indirect propagation of risk through network interactions (through $\mathbf{W}$). Note that by construction, $u_{i}=0$ for all pure obligees $i$, i.e. nodes with $\alpha_{i}=1$. This follows from the structure of $\mathbf{I}-\mathbf{A}$, where nodes corresponding to obligees contain only zero entries. Intuitively, pure obligees absorb risk but do not emit risk, and so their risk-based centrality measure is zero.
@@ 3.3 Impact of Risk Propagation on Mean Field
We close out this section with a central concern for contractor networks: the potential for risk amplification due to a contractor’s location within the network. While our model allows for flexible dynamics, it remains unclear how individual risk exposures respond to changes in the network structure. In particular, we seek to understand whether, for each $i\in\mathcal{V}$, a firm’s limiting failure probability $m_{i}$ is amplified (larger than $r_{i}$) when it interacts with riskier principals. To facilitate this analysis, we introduce a structural assumption on contracting organizations, which posits that contractors engage with organizations with higher risk. While this may not hold universally, we show that it is a sufficient condition for $\mathbf{m}\geq\mathbf{r}$ across the entire network.
@@
We assume for all intermediaries $i$ $\in\mathcal{V}$, the average risk of their principals is greater than or equal to their own inherent idiosyncratic risk score, i.e.
@@
While Section 3.3 may appear restrictive, it aligns with patterns observed in real-world contractor networks (bci2021supplychain). In practice, principals that subcontract work are often larger and hence less risky, whereas the subcontractors they engage with are typically smaller and more exposed to risk, which is in part due to a less rigorous vetting process (dietz2018mitigating). Although this assumption does not hold universally in our empirical setting in Section 5, we demonstrate that similar results emerge in our application.
@@
Under the opposite of Section 3.3, all inequalities hold in the reverse direction.
@@
The monotonicity in Section 3.3 can be viewed through the lens of potential theory: under Section 3.3 the mean failure probabilities evolve as a subharmonic function on the contractor graph, so the limiting vector $\mathbf{m}$ is the smallest subharmonic majorant of the initial risks $\mathbf{r}$ (chung1997spectral). Reversing the assumption yields the superharmonic analogue.
The first part of Section 3.3 shows that, under Section 3.3, the mean failure probability for each node $m_{i}$ evolves monotonically over time. Consequently, the expected global average loss $\mathbb{E}\mathopen{}\mathclose{{\left[\mathcal{L}(\mathbf{X}^{t})}}\right]$ is also monotone with respect to $t$ and converges to the global average risk associated with the stationary process. The intuition is straightforward: risk propagates through the network because intermediaries rely on principals who, on average, are at least as risky as themselves. Over time, this accumulation amplifies failure probabilities throughout the network. A central insight here is that under Section 3.3, incorporating network structure consistently yields a higher expected global loss than models that assume independent failures. The second part of Section 3.3 examines monotonicity with respect to the network-associated failure propagation rates $\alpha_{i}$ for all $i\in\mathcal{V}$. This highlights the robustness of Section 3.3: its conclusions hold across a wide range of $\alpha$ values and underscore how systemic risk is shaped jointly by idiosyncratic contractor risk and network position.
Our next assumption strengthens Section 3.3 by requiring that a contractor’s neighbors, on average, have not just a higher inherent risk but a risk level that exceeds the contractor’s by a fixed margin. We will then show the propagation of the margin in terms of its downstream cumulative risk measures. {assumption} We assume for all intermediaries $i$, the average risk of their neighbors is strictly greater than their own inherent idiosyncratic risk score, i.e. there exists a $\delta>0$ such that
@@ 4 Asymptotic Behavior of Stochastic Risk Process
The previous section analyzed the asymptotic behavior of the marginal average failure probabilities $m_{i}$ for all $i\in\mathcal{V}$. However, this approach ignores details about the network structure and the main feature that distinguishes this model from independent failure models: the potential for the limiting behavior of nodes to be correlated. In this section, we turn our attention to the full stochastic process $\mathbf{X}^{t}$. We first establish that $\mathbf{X}^{t}$ is a Markov chain that converges to a unique stationary distribution (proof deferred to Section 9), and then we quantify the rate of convergence to its stationary distribution in terms of the mixing time. We close out with a discussion on conditions under which the distribution of $\mathcal{L}(\mathbf{X}^{t})$ is stochastically dominated (a distributional extension of monotonicity) over $t$, highlighting the positive impact of network effects on tail risk.
@@ 4.1 Mixing Time of the Stochastic Process
Recall in Section 3.1 we showed that the convergence rate of the mean failure probabilities in our stochastic model depends on the rate of decay of $\mathbf{A}\mathbf{W}$. Here, we extend this insight to the full stochastic process, showing that the distribution over network failures also mixes depending on the rate of decay of $\mathbf{A}\mathbf{W}$. These results are, in a sense, quite surprising. Classical Markov chain arguments that operate directly on our state space $\{0,1\}^{n}$ would give bounds on the mixing time that scale with the size of the state space, i.e. exponentially in $n$ (resnick2013adventures). Such bounds are useless for the networks of interest here. Our analysis instead shows that the mixing time actually increases at a logarithmic rate in $n$ (for fixed accuracy $\epsilon$). Moreover, in the special but important case in which the contractor graph $G$ is a directed acyclic graph (DAG), the chain mixes in a finite number of steps that equals the depth of the DAG. Moreover, for DAGs the stationary distribution can be written in closed form by propagating probabilities along a topological order; see Fig. 5.
@@
As a result, if $G$ is a directed acyclic graph and $d>0$ denotes its maximum path length, then for all $\epsilon>0$ we have $t_{\textsf{mix}}(\epsilon)\leq d$. Similarly, if $G$ is a general graph then
@@
Now we note that if $G$ is a directed acyclic graph, then for any $t\geq d$ we have $(\mathbf{A}\mathbf{W})^{t}=0$. Hence, we get that $t_{\textsf{mix}}(\epsilon)\leq d$ for any $\epsilon>0$. For the general case, we use the fact that $\lVert(\mathbf{A}\mathbf{W})^{2}\rVert<1$ via Fig. 3. Plugging this into the above bound yields:
@@
We close out our discussion here with the case when the graph is a DAG. Because there are no directed cycles, the linear operator $(\mathbf{A}\mathbf{W})$ is nilpotent with index at most the depth $d$, and therefore the coupled chains coalesce after at most $d$ steps. This immediately implies that $t_{\textsf{mix}}(\epsilon)\leq d$ for all $\epsilon>0$, as described in Section 4.1. In the same spirit, the stationary distribution is computable in closed form by propagating probabilities layer by layer along a (reverse) topological ordering of the contractor graph $G$.

### v2

@@ 2 Network Model Definition
This example illustrates how the failure of a pure principal (e.g., $A$ defaults) can propagate through the network in the stochastic process. If $C$ cannot complete its contractual obligations to $E$ because $A$ fails, then $E$ may incur losses, even though it never directly contracted with $A$. Such indirect dependencies are not captured in standard models assuming independent risk, but they are central in our network-aware framework. This further underscores the role of intermediaries like $C$ in amplifying risk. Even with moderate idiosyncratic risk levels for $A$ and $B$, the dependency structure means that failures can cascade through the network, elevating systemic risk beyond what node-level scores would suggest.
@@ 2.1 Discussion of Modeling Assumptions
We treat each contractor $i$’s idiosyncratic risk score $r_{i}$ and network sensitivity parameter $\alpha_{i}$ as exogenously specified inputs to the model. In practice, these parameters are inferred from financial health indicators or historical default data, and likely take into account mild network risk indicators (Kim et al. 2019). However, regardless of how $r_{i}$ are estimated, our model allows for the direct incorporation of network effects on risk in contractor networks.
@@
Our stochastic process assumes that a contractor’s risk of failure increases linearly based on the impact of their neighbors via $w_{ij}$ (see Eq. 1). This additive structure simplifies both analysis and simulation, but it may fail to capture important nonlinearities in real-world contagion effects. For instance, a contractor may be robust to isolated failures but vulnerable to risks beyond a certain threshold, such as in the threshold contagion model like Watts (2002).
@@
Our work assumes that network-induced risk depends on the proportion of a contractor’s total subcontracted value attributed to each subcontractor, as encoded by the normalized edge weights $w_{ij}$. This formulation reflects a reasonable first-order approximation where risk exposure grows with financial dependence on risky neighbors. Although real-world contracts vary in risk beyond their dollar value, introducing such heterogeneity would significantly complicate the model without materially changing our core theoretical insights.
@@ 3 Mean Field Analysis of Expected Risk
We start off our analysis by considering the marginal expected failure probabilities of the stochastic process $\mathbf{X}^{t}$ for each node $i$. We will later see that this corresponds to a modified Friedkin-Johnsen model in the opinion dynamics literature (Friedkin and Johnsen 1990), and calculate a closed-form expression for the mean failure probabilities. We also exploit this representation to describe an eigenvector-based centrality measure, assigning scores to each node in the graph corresponding to their risk-based centrality within the contractor network. We close this section by providing a simple monotonicity condition under which the mean failure probabilities for each node increase due to network effects.
@@ 3.1 Expected Failure Probabilities
The operator $(\mathbf{I}-\mathbf{A}\mathbf{W})^{-1}$ admits multiple interpretations across related literatures. In Markov chain theory, it parallels the fundamental matrix of absorbing chains, where each entry gives the expected number of visits to a state prior to absorption (Kemeny et al. 1969). From a graph-theoretic perspective, it plays a role similar to the pseudoinverse of the graph Laplacian, which encodes mean hitting times, commute times, and effective resistances (Lovász 1993, Doyle and Snell 1984, Chung 1997). In our context, this highlights that $(\mathbf{I}-\mathbf{A}\mathbf{W})^{-1}$ aggregates contributions from all higher-order paths in the contractor network, providing an interpretable link between local idiosyncratic risks and their amplified global effects.
@@
lemmaSecondNormBounded Let $\mathbf{A}$ be a diagonal matrix with $\mathbf{A}_{ii}=\alpha_{i}\text{ for all }i\in\mathcal{V}$ and let $\mathbf{W}$ be the row-substochastic adjacency matrix of the contractor graph $G$. Then we have $\lVert(\mathbf{A}\mathbf{W})^{2}\rVert<1$ and $(\mathbf{A}\mathbf{W})^{t}\to 0$ as $t\to\infty$.
This bound is useful for showing that the Neumann series of $\mathbf{A}\mathbf{W}$ converges, which we formalize in the following corollary: {restatable}corollaryNeumann Let $\mathbf{A}$ be a diagonal matrix with $\mathbf{A}_{ii}=\alpha_{i}\text{ for all }i\in\mathcal{V}$ and let $\mathbf{W}$ be the row-substochastic adjacency matrix of the contractor graph $G$. Then $(\mathbf{I}-\mathbf{A}\mathbf{W})$ is invertible, and the Neumann series $\sum_{t=0}^{\infty}(\mathbf{A}\mathbf{W})^{t}$ converges to $(\mathbf{I}-\mathbf{A}\mathbf{W})^{{-1}}$.
@@
Relation to Friedkin and Johnsen (1990).
Our model is closely related to the social influence model under static conditions introduced by Friedkin and Johnsen (1990), which describes a process in which an individual’s opinion at time $t+1$ is represented as a real-valued linear function of both their innate opinion (determined by exogenous variables) and the opinions of their in-neighbors in the previous time step $t$. Let $\mathbf{m}^{t}$ denote the vector of opinions at time $t\geq 0$, where $\mathbf{r}=\mathbf{m}^{0}$ denotes the vector of inherent opinions that each individual holds at time $t=0$. Then opinions evolve according to the following dynamics for all $t\in\mathbb{N}$:
@@
Adapting their model to our context of risk, we interpret a node’s opinion as its probability of failure, which is dependent on failure probabilities of subcontractors according to the transition matrix $\mathbf{W}$ in the same manner that individuals in the social influence model take their neighbors’ opinions into account. Thus in our context we require that “opinions” take values in $[0,1]$. Moreover, to ensure that $\mathbf{m}^{t}$ maintains its interpretation as a vector of probabilities for all $t\in\mathbb{N}$, we also require $\mathbf{W}$ to be row-substochastic and that $\alpha,\gamma>0$ satisfy $\alpha+\gamma=1$. Then we can rewrite Equation 7 as $\mathbf{m}^{t+1}=(1-\alpha)\mathbf{r}+\alpha\mathbf{W}\mathbf{m}^{t}.$
Comparing with our dynamics in Equation 3, the main difference from Friedkin and Johnsen (1990) is that they assume a common bias parameter $\alpha$, whereas our model allows node-specific bias parameters by replacing $\alpha$ and $\gamma$ with diagonal matrices $\mathbf{A}$ and $(\mathbf{I}-\mathbf{A})$. Note that in contrast to the Friedkin-Johnsen model, we consider a directed network, so to handle nodes $i\in\mathcal{V}$ that lack in- or out-edges we fix their $\alpha_{i}$ as described in Section 2. More broadly, their framework models the deterministic evolution of average opinions, while ours defines a discrete stochastic process over binary outcomes, yielding the full distribution of failures across the network rather than only mean behavior.
@@ 3.2 Identifying the Impact of Network Structure
where $n$ is the total number of nodes in the network. This expression shows that the overall failure risk can be rewritten as a network-adjusted re-weighting of the individual idiosyncratic risk scores $\mathbf{r}$. The term $(\mathbf{I}-\mathbf{A}\mathbf{W})^{-1}$ highlights that risk exposure is not limited to direct neighbors since it aggregates contributions from all paths in the network, with longer paths down-weighted by successive products of exposure probabilities.
@@
This definition is analogous to traditional PageRank in web search algorithms (Brin and Page 1998, Page et al. 1999, Kleinberg 1999), where importance is assigned based on structural connectivity. Here, however, the risk-based centrality captures how individual nodes influence the system-wide failure probability. It accounts for both direct contributions from individual risk levels (through $\mathbf{A}$) and the indirect propagation of risk through network interactions (through $\mathbf{W}$). Note that by construction, $u_{i}=0$ for all pure obligees $i$, i.e. nodes with $\alpha_{i}=1$. This follows from the structure of $\mathbf{I}-\mathbf{A}$, where nodes corresponding to obligees contain only zero entries. Intuitively, pure obligees absorb risk but do not emit risk, and so their risk-based centrality measure is zero.
@@ 3.3 Impact of Risk Propagation on Mean Field
We close out this section with a central concern for contractor networks, the potential for risk amplification due to a contractor’s location within the network. While our model allows for flexible dynamics, it remains unclear how individual risk exposures respond to changes in the network structure. In particular, we seek to understand whether, for each $i\in\mathcal{V}$, a firm’s limiting failure probability $m_{i}$ is amplified (larger than $r_{i}$) when it interacts with riskier principals. To facilitate this analysis, we introduce a structural assumption on contracting organizations, which posits that contractors engage with organizations with higher risk. While this may not hold universally, we show that it is a sufficient condition for $\mathbf{m}\geq\mathbf{r}$ across the entire network.
@@
For all intermediaries $i$ $\in\mathcal{V}$, the average risk of their principals is greater than or equal to their own inherent idiosyncratic risk score, i.e.
@@
While Section 3.3 may appear restrictive, it aligns with patterns observed in real-world contractor networks (Business Continuity Institute 2021). In practice, principals that subcontract work are often larger and hence less risky, whereas the subcontractors they engage with are typically smaller and more exposed to risk, which is in part due to a less rigorous vetting process (Dietz 2018). Although this assumption does not hold universally in our empirical setting in Section 5, we demonstrate that similar results emerge in our application.
@@
Under the opposite of Section 3.3, all inequalities hold in the reverse direction for principals and intermediaries.
@@
The monotonicity in Section 3.3 can be viewed through the lens of potential theory. Under Section 3.3 the mean failure probabilities evolve as a subharmonic function on the contractor graph, so the limiting vector $\mathbf{m}$ is the smallest subharmonic majorant of the initial risks $\mathbf{r}$ (Chung 1997). Reversing the assumption yields the superharmonic analogue.
The first part of Section 3.3 shows that, under Section 3.3, the mean failure probability for each node $m_{i}$ evolves monotonically over time. Consequently, the expected global average loss $\mathbb{E}\mathopen{}\mathclose{{\left[\mathcal{L}(\mathbf{X}^{t})}}\right]$ is also monotone with respect to $t$ and converges to the global average risk associated with the stationary process. The intuition is straightforward: risk propagates through the network because intermediaries rely on principals who, on average, are at least as risky as themselves. Over time, this accumulation amplifies failure probabilities throughout the network. A central insight here is that under Section 3.3, incorporating network structure consistently yields a higher expected global loss than models that assume independent failures. The second part of Section 3.3 examines monotonicity with respect to the network-associated failure propagation rates $\alpha_{i}$ for all $i\in\mathcal{V}$. This highlights the robustness of Section 3.3, since its conclusions hold across a wide range of $\alpha$ values and underscore how systemic risk is shaped jointly by idiosyncratic contractor risk and network position.
Our next assumption strengthens Section 3.3 by requiring that a contractor’s neighbors, on average, have not just a higher inherent risk but a risk level that exceeds the contractor’s by a fixed margin. We will then show the propagation of the margin in terms of its downstream cumulative risk measures. {assumption} For all intermediaries $i$, the average risk of their neighbors is strictly greater than their own inherent idiosyncratic risk score, i.e. there exists a $\delta>0$ such that
@@ 4 Asymptotic Behavior of Stochastic Risk Process
The previous section analyzed the asymptotic behavior of the marginal average failure probabilities $m_{i}$ for all $i\in\mathcal{V}$. However, this approach ignores details about the network structure and the main feature that distinguishes this model from independent failure models, the potential for the limiting behavior of nodes to be correlated. In this section, we turn our attention to the full stochastic process $\mathbf{X}^{t}$. We first establish that $\mathbf{X}^{t}$ is a Markov chain that converges to a unique stationary distribution (proof deferred to Section 9), and then we quantify the rate of convergence to its stationary distribution in terms of the mixing time.
@@ 4.1 Mixing Time of the Stochastic Process
Recall in Section 3.1 we showed that the convergence rate of the mean failure probabilities in our stochastic model depends on the rate of decay of $\mathbf{A}\mathbf{W}$. Here, we extend this insight to the full stochastic process, showing that the distribution over network failures also mixes depending on the rate of decay of $\mathbf{A}\mathbf{W}$. These results are, in a sense, quite surprising. Classical Markov chain arguments that operate directly on our state space $\{0,1\}^{n}$ would give bounds on the mixing time that scale with the size of the state space, i.e. exponentially in $n$ (Resnick 2013). Such bounds are useless for the networks of interest here. Our analysis instead shows that the mixing time actually increases at a logarithmic rate in $n$ (for fixed accuracy $\epsilon$). Moreover, in the special but important case in which the contractor graph $G$ is a directed acyclic graph (DAG), the chain mixes in a finite number of steps that equals the depth of the DAG. Moreover, for DAGs the stationary distribution can be written in closed form by propagating probabilities along a topological order; see Fig. 5.
@@
As a result, if $G$ is a directed acyclic graph and $d>0$ denotes its maximum path length, then for all $\epsilon>0$ we have $t_{\textsf{mix}}(\epsilon)\leq d+1$. Similarly, if $G$ is a general graph then
@@
Now we note that if $G$ is a directed acyclic graph, then for any $t>d$ we have $(\mathbf{A}\mathbf{W})^{t}=0$. Hence, we get that $t_{\textsf{mix}}(\epsilon)\leq d+1$ for any $\epsilon>0$. For the general case, we use the fact that $\lVert(\mathbf{A}\mathbf{W})^{2}\rVert<1$ via Fig. 3. Plugging this into the above bound yields:
@@
We close out our discussion here with the case when the graph is a DAG. Because there are no directed cycles, the linear operator $(\mathbf{A}\mathbf{W})$ is nilpotent with index at most the depth $d+1$, and therefore the coupled chains coalesce after at most $d+1$ steps. This immediately implies that $t_{\textsf{mix}}(\epsilon)\leq d+1$ for all $\epsilon>0$, as described in Section 4.1. In the same spirit, the stationary distribution is computable in closed form by propagating probabilities layer by layer along a (reverse) topological ordering of the contractor graph $G$.